ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 160, 2007

Bagno, E. and R. Biagioli Colored-descent representations of complex reflection groups G(r, p, n)

317

Berele, A. Approximate values of the multiplicities in the arm of the cocharacter sequence of M_{2,1}

371

Biagioli, R. See E. Bagno

Buckner, J. and M. Dugas Quasi-localizations of ℤ  349

Dinur, I., E. Friedgut, G. Kindler and R. O’Donnell On the Fourier tails of bounded functions over the discrete cube

389

Dugas, M. See J. Buckner

Friedgut, E. See I. Dinur

Helm, D. On maps between modular Jacobians and Jacobians of Shimura curves

61

Hirayama, M. and Y. Pesin Non-absolutely continuous foliations  173

Kindler, G. See I. Dinur

Kowalski, E. Exponential sums over definable subsets of finite fields 219

Ledrappier, F. and O. Sarig Invariant measures for the horocycle flow on periodic hyperbolic surfaces

281

Magazanik, E. and M. A. Perles Relatively convex subsets of simply connected planar sets

143

Magazanik, E. and M. A. Perles Generalized convex kernels of simply connected L_{n} sets in the plane

157

Mattarei, S. The orders of nonsingular derivations of Lie algebras of characteristic two

23

Merimovich, C. Prikry on extenders, revisited  253

O’Donnell, R. See I. Dinur

Perles, M. A. See E. Magazanik

Pesin, Y. See M. Hirayama

Phillips, N. C. Examples of different minimal diffeomorphisms giving the same C*-algebras 189 Rubinstein, J. and I. Shafrir The distance between homotopy

classes of S^{1}-valued maps in multiply connected domains

41

Sarig, O. See F. Ledrappier Shafrir, I. See J. Rubinstein Shilon, O. and B. Weiss Universal minimal topological dynamical

systems

119

Weiss B. See O. Shilon

Xi, L.-F. Quasi-Lipschitz equivalence of fractals  1